One cellular phone carrier charges $26.50 a month plus $.15 a minute for local calls. Another carrier charges $14.50 a month and $.25 a minute for local calls. For how many minutes is the cost of the plans the same?
Answer by Boreal(15235)

Set them equal: 26.50+0.15x=14.50+0.25x, where x= minutes
12=0.10x
x=120 minutes
check
$44.50=$44.50